### v3

# Low-Resource Dialect Adaptation of Large Language Models: A French Dialect Case-Study

Abstract
Despite the widespread adoption of Large Language Models (LLMs), their strongest capabilities remain largely confined to a small number of high-resource languages for which there is abundant training data. Recently, continual pre-training (CPT) has emerged as a means to fine-tune these models to low-resource regional dialects. In this paper, we study the use of CPT for dialect learning under tight data and compute budgets. Using low-rank adaptation (LoRA) and compute-efficient continual pre-training, we adapt three LLMs to the Québec French dialect using a very small dataset and benchmark them on the COLE suite. Our experiments demonstrate an improvement on the minority dialect benchmarks with minimal regression on the prestige language benchmarks with around 1% of model parameters updated. Analysis of the results demonstrate that gains are highly contingent on corpus composition. These findings indicate that CPT with parameter-efficient fine-tuning (PEFT) can narrow the dialect gap by providing cost-effective and sustainable language resource creation, expanding high-quality LLM access to minority linguistic communities. To support reproducibility and broaden access, we release the first Québec French LLMs on Hugging Face.
Keywords: Continual pre-training, Dialect adaptation, Parameter-efficient fine-tuning, LoRA, Low-resource languages, Québec French, Language equity, Domain adaptation
[…]
Low-Resource Dialect Adaptation of Large Language Models:
A French Dialect Case-Study
Eeham Khan1, Firas Saidani2, Owen Van Esbroeck1, Richard Khoury2, Leila Kosseim1
1 Computational Linguistics at Concordia (CLaC) Laboratory
Department of Computer Science and Software Engineering
Concordia University, Montréal, Québec, Canada
{eeham.khan, leila.kosseim}@concordia.ca, owen.vanesbroeck@mail.concordia.ca
2 Group for Research in Artificial Intelligence of Laval University (GRAIL)
Department of Computer Science and Software Engineering
Université Laval, Québec, Canada
{richard.khoury, firas-mustapha.saidani.1}@ulaval.ca
Abstract content
1.  Introduction
In recent years, Large Language Models (LLMs) have emerged as powerful and versatile tools for many applications, driving progress in tasks such as text summarization, text and code generation, and open-domain dialogue systems. These LLMs are pretrained on extensive corpora to maximize general-purpose performance across a wide variety of downstream tasks.
Despite their success, most widely-used LLMs are trained on predominantly English datasets. When other languages are included in multilingual training datasets, they are collected from the high-resource prestige dialect of the language and have little coverage of low-resource regional dialects. As a result, they often struggle with local vocabulary, orthographic variants, idiomatic expressions, and code-switching phenomena that are commonly found in regional dialects. This issue, coined the dialect gap Kantharuban et al. (2023), limits their effectiveness for millions of users of these minority language varieties, thereby perpetuating inequities in access to artificial intelligence technologies.
One possible approach to addressing the dialect gap is to train LLMs using regional dialectal data. However, full model training is prohibitively expensive, while conventional fine-tuning approaches can be inefficient or lead to overfitting on relatively small dialectal datasets. Continual Pre-training (CPT) on unlabeled data offers a practical compromise: by exposing models to large volumes of dialect-specific text, CPT enhances regional linguistic coverage without fully discarding the general knowledge encoded during initial pre-training (Gururangan et al., 2020; Sarkar et al., 2022; Lee et al., 2020).
In this paper, we explore the ability and limitations of CPT to adapt LLMs to a regional dialect using a small amount of unlabeled regional texts. As a case study, we consider specifically the case of the Québec regional dialect of French, also called Québécois. To make adaptation feasible with less compute resources, we employ low-rank adaptation (LoRA) and gradient checkpointing as parameter-efficient strategies. The adapted models are then benchmarked on a subset of the COLE suite of French tasks Beauchemin et al. (2025a), including both Québec French and prestige French benchmarks, to assess their ability to balance dialectal adaptation while retaining general competence in French.
Our key contributions in this paper are:
  1. 1.
    We demonstrate a compute-efficient CPT pipeline (LoRA + gradient checkpointing) that updates $\leq\!1\%$ of parameters while sustaining performance on modest hardware (see Section 4). We show that it is possible to adapt an LLM to a regional dialect using only a very small corpus of 86M tokens in our experiments, which is significantly below typical CPT data budgets.
  2. 2.
    Using this methodology, we train and release the first open-weight LLMs adapted specifically to Québec French, and evaluate them on a subset of the COLE benchmark suite. These tests highlight the capacity and limitations of CPT dialectal training (Section 6).
  3. 3.
    We provide complete training configurations, data-processing scripts, and evaluation pipelines for direct transfer to other dialects and low-resource varieties on GitHub [^1] https://github.com/CLaC-Lab/QuebecLLM-CPT. We also release our Québec French LLMs on HuggingFace [^2] https://huggingface.co/QuebecLLM.
2.  Background and Related Work
2.1.  Continual Pre-training
Continual Pre-training continues training a model for its original self-supervised objective (e.g., masked or causal language modeling) using an out-of-domain corpus. The objective is to extend the model’s capabilities using additional, unlabeled datasets. CPT differs from supervised fine-tuning, which uses smaller task-specific labeled datasets.
Although CPT is related to continual learning (CL), it is more appropriate for practical specialization rather than lifelong multi-task learning. CL methods such as regularization (Kirkpatrick et al., 2017; Aljundi et al., 2018), replay (Lopez-Paz and Ranzato, 2017; Chaudhry et al., 2019), and gradient orthogonalization (Farajtabar et al., 2020) mitigate catastrophic forgetting across many tasks. On the other hand, CPT typically adapts to a single new distribution without explicit replay or regularizers, retaining the model’s original capability through careful optimization and small learning rates (Gururangan et al., 2020).
While full-parameter fine-tuning is often prohibitive for LLMs due to compute time and cost, low-rank adaptation (LoRA) (Hu et al., 2022) allows updates to only a small fraction of model parameters while preserving pretrained weights. Recent advances combine LoRA with quantization-aware training (e.g., L4Q (Jeon et al., 2025)) to further reduce costs. Other parameter-efficient fine-tuning (PEFT) techniques such as adapters (Houlsby et al., 2019; Pfeiffer et al., 2021), prefix/prompt tuning (Li and Liang, 2021; Lester et al., 2021), BitFit (Ben Zaken et al., 2022), and learned attention/FFN scaling (Liu et al., 2022) reduce memory and computation costs while achieving competitive performances. For resource-constrained continual pre-training, LoRA and related methods can be combined with mixed precision, gradient checkpointing, and quantization to keep adaptation feasible (Dettmers et al., 2023).
$2.$  Dialect Adaptation
Adapting pretrained LLMs to new linguistic varieties can be framed as domain-adaptive pre-training (DAPT) or language-adaptive pre-training (LAPT), where models are exposed to additional unlabeled target-distribution text (Gururangan et al., 2020; Howard and Ruder, 2018; Chronopoulou et al., 2019). This improves coverage of distribution-specific lexical items, morphology, and style while preserving general knowledge of the base model. Prior work has applied this to new domains (e.g. biomedical or legal) Lee et al. (2020); Chalkidis et al. (2020) and low-resource languages (Beltagy et al., 2019; Koto et al., 2025; Mansha, 2025).
CPT has become a dominant approach for adapting LLMs to low-resource languages and regional dialects Elhady et al. (2025).
Such systems have been developed for a variety of languages and dialects, trained on datasets as small as 350M tokens. Recent work also shows that mixing English with the target language during CPT can be critical for preserving downstream abilities Elhady et al. (2025).
Language Tokens Reference
Québécois 0.085B Ours
Galician (Carballo) 0.340B  Proxecto NOS (2024)
Catalan (CataLlama) 0.445B  CataLlama Team (2024)
Sorbian 1.2B  Fishel et al. (2025)
Basque (Latxa) 4.2B  Etxaniz et al. (2024)
Kazakh (Sherkala-Chat) 45B (Koto et al., 2025)
German (LeoLM) 65B  Keller et al. (2023)
Hindi (Nemotron-Mini) 400B  Dabre et al. (2025)
Vietnamese (VinaLlama) 500B  VietAI Research (2023)
Arabic (ALLaM) 600B to 1200B  Bari et al. (2025)
Table 1: Regional languages for which LLMs have been trained using CPT, along with the number of training tokens.
French LLMs such as CamemBERT and FlauBERT established strong monolingual baselines (Martin et al., 2020; Le et al., 2020), while multilingual encoders (e.g., XLM-RoBERTa) remain competitive for French (Conneau et al., 2020). However, these are trained on the prestige dialect of French used in France. The Québec dialect of French, or Québécois, differs from the prestige dialect in orthography, phonology-to-orthography conventions, anglicisms, idioms, and code-switching patterns. In addition, Québécois resources are scarce, compared to prestige French or to other languages listed in Table 1. These characteristics make CPT a viable option for dialectal adaptation.
3.  Datasets
3.1.  Data sources
To adapt models to Québécois, we collected a corpus of documents from a variety of Québec French sources spanning news, blogs, transcribed speech, and social media. All copyrighted materials were obtained and used with the explicit permission of the rights holders.
Source Type Years Size License
1. BEQ ebooks Books 1800–1960 15.44M Public domain
2. Wikipedia (QC category) Articles 2008–2025 31.15M CC-BY-SA 3.0
3. CN2i – Le Soleil News articles 2000–2025 5.86M Copyrighted
4. CRIFUQ oral transcripts Interviews 2012–2019 0.76M Non-commercial
5. Facebook Le Soleil Comments 2020–2023 2.40M ToS-restricted
6. Depotoir.ca Forum posts 2009–2025 22.82M Non-commercial
7. MontrealRacing.com Forum posts 2003–2025 4.23M Non-commercial
8. YouTube (QC channels) Comments 2006–2025 2.15M Non-commercial
9. Reddit (QC communities) Comments 2025 1.76M Non-commercial
Total 86.57M
Table 2: Sources of the Québec French corpus collected and used for CPT. Sizes are token counts using the CroissantLLM tokenizer.
The data sources we used are reported in Table 2. We can distinguish two categories: sources that include formal texts (numbers 1 to 3 in Table 2) which make up 60% (52.45M tokens) of our corpus, and those that include informal (spoken or user-generated) texts (numbers 4 to 9) which make up the remaining 40% (34.12M tokens). These sources are described below:
  1. 1.
    BEQ Ebooks (15.44M tokens):
    Public-domain literary texts in Québec French, sourced from the Bibliothèque électronique du Québec (BEQ). These are primarily literary texts written in normative Québec French. This source offers diachronic coverage of historical Québecois literary language, enriching the overall corpus with stylistic and temporal diversity.
  2. 2.
    Wikipedia (QC) (31.15M tokens):
    Articles retrieved from the Québec portal of Wikipedia, covering Québec topics and presumed to be primarily written by Québécois people in formal prose. This source contributes broad, topic-specific coverage written in Québec French.
  3. 3.
    CN2i – Le Soleil Newspaper (5.86M tokens):
    News articles from the Québec City newspaper, Le Soleil, written in formal journalistic prose. This source offers high-quality formal written texts covering a variety of current events.
  4. 4.
    CRIFUQ Oral Transcripts (0.76M tokens):
    Interview transcriptions from the Centre de recherche interuniversitaire sur le français en usage au Québec (CRIFUQ). It is our only collection of spontaneous speech in informal Québec French. It thus provides examples of features typical of Québec French speech (e.g., ellipsis, disfluencies, phonology-to-orthography variation).
  5. 5.
    Facebook Comments – Le Soleil (2.40M tokens):
    User comments on posts by the newspaper Le Soleil, collected via the Facebook Graph API. These are examples of user-generated informal public discourse by newspaper-reading (thus older and more educated) individuals. It captures unedited conversational patterns and informal register.
  6. 6.
    Depotoir.ca (22.82M tokens):
    Public posts on the Québec-based forum Depotoir.ca. These posts contain informal, colloquial, and regionally-specific Québécois, including slang and argot. This source contributes content in vernacular Québec French rich in sociolinguistic variations, including non-standard orthography and expressive discourse, and covering a variety of cultural, social, and general discussion topics.
  7. 7.
    MontrealRacing.com (4.23M tokens):
    Public posts from a Montreal-based automotive enthusiast forum. This source provides other examples of informal Québec French, this time rich with technical terminology and slang.
  8. 8.
    YouTube Comments (2.15M tokens):
    Public user comments from Québec YouTube channels, collected via the YouTube API. These are texts of informal conversational Québécois, often containing regionalisms, and code-switching. They provide examples of more youthful writing patterns and slang.
  9. 9.
    Reddit – Québec Communities (1.76M tokens):
    Public comments from francophone Québec subreddits, collected via Reddit API. Another example of informal, colloquial Québec French, including slang and regional features. Unlike the other comments, this source provides examples of more sustained informal conversations.
Overall, our Québec French corpus spans 86.57M tokens, which is significantly below typical CPT data budgets used for dialectal/low-resource adaptations, as shown in Table 1.
3.2.  Data Pre-Processing
To use the Québec French corpus for CPT, we applied light pre-processing to preserve dialectal markers and standardize formatting, allowing it to be used as a single, consistent dataset. To that end, each line of text was stripped of byte-order markers, HTML or Wiki tags, and extra whitespaces. Empty lines were removed, but paragraph boundaries and sentence-level markers were preserved as much as possible. Only trivial corrections, such as collapsing repeated spaces, were applied, ensuring that distinct orthographic patterns were preserved.
4.  Training
To reduce the cost of CPT, we employed LoRA (Hu et al., 2022) instead of updating all model parameters. Following common practice, we targeted the attention projections ($q$, $k$, $v$, $o$) and feed-forward layers (up, down, gate) with rank $r=16$, $\alpha=32$, and dropout of $0.1$. Trainable parameters are cast to float32, while frozen parameters remain in float16. Additionally, we enabled gradient checkpointing to further lower memory requirements. This strategy yielded a trainable parameter ratio of around $1\%$ of the full model, making dialectal adaptation feasible on modest hardware.
We performed CPT with a causal language modeling (CLM) objective. The model was exposed to our Québec French corpus (see Section 3) in a single pass without domain interleaving or replay of the previously used prestige French training data. Due to computational constraints, sequences were limited to $1024$ tokens with a stride of $512$, producing overlapping chunks that preserve context for long documents. We trained each model for 3 and 6 epochs over the corpus using the AdamW optimizer with a weight decay of $0.01$. The learning rate was set to $1\times 10^{-5}$ with cosine decay and a warm-up ratio of $0.1$. To avoid numerical instability during the training process, we clipped gradients at a norm of $1.0$.
The effective batch size in sequences can be described as:
$\text{EffBatch}_{\text{seq}}=b\times a\times d,$
where $b$ is the number of sequences per device (micro-batch), $a$ is the gradient-accumulation steps, and $d$ is the number of devices. With a sequence length of $1024$ tokens, the effective batch size in tokens was therefore $\text{EffBatch}_{\text{tok}}=\text{EffBatch}_{\text{seq}}\times 1024$. See Appendix A for more details.
5.  Experimental Setup
5.1.  Base Models
We developed the Québec French models by conducting CPT on the following base models:
  • CroissantLLMChat-v0.1 (1.35B): a bilingual (English/French) open-source model. This serves as a strong dialect-aware baseline that is already exposed to French text.
  • Llama-3.2-1B: a lightweight general-purpose model included as a baseline for multilingual models.
  • Llama-3.1-8B: a high-capacity general-purpose model included as a stronger large-scale multilingual baseline for comparison.
All the models were evaluated both in their base forms and after 3 then 6 epochs of CPT following the training regime described in Section 4.
5.2.  Evaluation Benchmarks
We evaluated our models on 8 tasks of the COLE French-language benchmark Beauchemin et al. (2025a). To evaluate both the models’ acquisition of Québec French and their retention of general abilities after CPT, we chose 4 Québec French tasks (QFrCoLA, QFrBLiMP, QFrCoRE and QFrCoRT) and 4 prestige French tasks (AlloCiné, PAWS-X, Fr-BoolQ, and MMS). These 8 tasks are described below:
  1. 1.
    QFrCoLA Beauchemin and Khoury (2025): A grammatical acceptability task. The dataset consists of a total of 25,153 individual sentences classified as grammatical or ungrammatical. The sentences are drawn from a normative Québec French resource, and grouped by phenomena (syntax, morphology, semantics, anglicism). We used the 7,546 test samples for evaluation.
  2. 2.
    QFrBLiMP Beauchemin et al. (2025c): A grammatical acceptability task. The dataset is composed of 1,761 carefully-edited sentence pairs, one correct and the other with a common linguistic mistake. The task consists of identifying the correct sentence. Linguistic mistakes cover 20 attested phenomena from the “Banque de dépannage linguistique” (BDL), an official Québec government grammar source, annotated by native speakers. We used the 529 test samples for evaluation.
  3. 3.
    QFrCoRE Beauchemin et al. (2025b): A definition matching task. The dataset includes 4,633 Québec French multi-word expressions, with 10 possible definitions, and the correct one must be identified. The expressions and correct definitions were sourced from Québec regional-language collections. We used all 4,633 samples as a test set for evaluation.
  4. 4.
    QFrCoRT Beauchemin et al. (2025b): A definition matching task. This task is similar to QFrCoRE, but for 201 single-word Québec French idioms. We used all 201 samples as a test set for evaluation.
  5. 5.
    AlloCiné Blard (2020): A binary sentiment classification of 200,000 movie reviews in French. We used the test set of 20,000 samples for evaluation.
  6. 6.
    PAWS-X Yang et al. (2019): A paraphrase detection task. This task consists in identifying sentences that are paraphrases of each other. This multilingual dataset is composed of 23,659 professionally translated sentence pairs, but we used the test set of 2,000 French samples for evaluation.
  7. 7.
    Fr-BoolQ Clark et al. (2019): A binary reading comprehension task. This task is composed of 15,942 yes/no questions paired with passages containing the answer. This dataset is also multilingual, but we used the French subset of 178 test samples.
  8. 8.
    MMS Augustyniak et al. (2023): A sentiment analysis task. This task provides text snippets taken from 79 corpora from various domains, and requires identifying the sentiment in a binary- or trinary-class schema, depending on the corpus. This dataset is also multilingual, but we used the French subset of 63,190 test samples.
Table 3 summarizes all of these tasks with respect to their label distribution.
Québécois French
Label
QFrCoLA
QFrBLiMP
QFrCoRE
QFrCoRT
Allociné
Fr-BoolQ
MMS
PAWS-X
0 30.5 48.0 10.7 12.8 52.0 50.0 39.9 54.8
1 69.5 52.0 9.9 8.1 47.9 50.0 20.5 45.1
2 9.0 11.1 39.5
3 10.2 9.3
4 10.1 7.6
5 10.3 10.5
6 9.3 11.6
7 9.6 10.5
8 10.3 9.3
9 10.1 8.7
Table 3: Label frequency (%) across the selected COLE tasks.
COLE Québec-French Tasks
[…]
CroissantLLM family
Croissant (Base) $27.33$ $32.84$ $2.91$ $4.59$
Croissant (3 epochs CPT) $26.56$ 0.77-0.77 $32.84$ $0.00$ $2.77$ 0.14-0.14 $5.87$ +1.28+1.28 +0.09+0.09
Croissant (6 epochs CPT) $\boxed{46.22}$ +18.89+18.89 $\boxed{35.90}$ +3.06+3.06 $1.76$ 1.15-1.15 $1.61$ 2.98-2.98 +4.45+4.45
Llama-3.2-1B family
Llama-1B (Base) $45.84$ $32.44$ $1.83$ $2.00$
Llama-1B (3 epochs CPT) $27.89$ 17.95-17.95 $34.92$ +2.45+2.45 $3.68$ +1.85+1.85 $4.54$ +2.54+2.54 2.78-2.78
Llama-1B (6 epochs CPT) $42.45$ 3.39-3.39 $\boxed{35.90}$ +3.46+3.46 $4.58$ +2.75+2.75 $3.88$ +1.88+1.88 +1.18+1.18
Llama-3.1-8B family
Llama-8B (Base) $41.04$ $32.44$ $5.84$ $10.54$
Llama-8B (3 epochs CPT) $40.99$ 0.05-0.05 $32.44$ $0.00$ $8.12$ +2.28+2.28 $12.33$ +1.79+1.79 +1.01+1.01
Llama-8B (6 epochs CPT) $41.66$ +0.62+0.62 $32.44$ $0.00$ $\boxed{8.91}$ +3.07+3.07 $\boxed{13.73}$ +3.19+3.19 +1.72+1.72
Table 4: Results on COLE Québec French tasks only. $\boxed{Boxed}$ marks global best within each column. $\Delta$ columns show improvement over the corresponding base model; gray cells summarize the average $\Delta$ across Québec French tasks.
General French Tasks
[…]
CroissantLLM family
Croissant (Base) $36.18$ $40.57$ $\boxed{50.75}$ $19.37$
Croissant (3 epochs CPT) $41.10$ +4.92+4.92 $41.55$ +0.98+0.98 $44.30$ 6.45-6.45 $26.42$ +7.05+7.05 +1.63+1.63
Croissant (6 epochs CPT) $50.39$ +14.21+14.21 $\boxed{50.07}$ +9.50+9.50 $32.06$ 18.69-18.69 $28.51$ +9.14+9.14 +3.54+3.54
Llama-3.2-1B family
Llama-1B (Base) $\boxed{70.96}$ $35.42$ $50.29$ $19.11$
Llama-1B (3 epochs CPT) $45.05$ 25.91-25.91 $45.11$ +9.69+9.69 $40.80$ 9.49-9.49 $19.89$ +0.78+0.78 6.23-6.23
Llama-1B (6 epochs CPT) $49.77$ 21.19-21.19 $35.56$ +0.14+0.14 39.219.21 11.08-11.08 $19.12$ +0.01+0.01 8.03-8.03
Llama-3.1-8B family
Llama-8B (Base) $34.37$ $34.10$ $40.20$ $22.17$
Llama-8B (3 epochs CPT) $53.95$ +19.58+19.58 $47.89$ +13.79+13.79 $39.13$ 1.07-1.07 $30.82$ +8.65+8.65 +10.24+10.24
Llama-8B (6 epochs CPT) $57.11$ +22.74+22.74 $49.81$ +15.71+15.71 $38.17$ 2.03-2.03 37.82 +15.65+15.65 +13.02+13.02
Table 5: Results on general French tasks only. $\boxed{Boxed}$ marks global best within each column. $\Delta$ columns show improvement over the corresponding base model; gray cells summarize the average $\Delta$ across prestige French tasks.
None of the texts in the Québec French tasks (QFrCoLA, QFrBLiMP, QFrCoRE, QFrCoRT) were built from, or were included in the Québec French unlabeled corpus used for CPT (see Section 3.1).
6.  Results and Analysis
[…]
We analyze the results on three aspects. First, we consider the models’ ability to acquire Québécois (see Section 6.1). Second, we look at the models’ retention of prestige French abilities (see Section 6.2). Finally, we study the adaptation-retention trade-off of the models, or their ability to balance learning and remembering language (see Section 6.3).
6.1.  Québec French Acquisition
Figure 1 shows the perplexity of the three models before (epoch 0) and after each of the six epochs of CPT. To compute perplexity scores, we used 5% of the COLE test sets during training. Because this is training perplexity (with overlap), it likely underestimates true generalization error. We therefore treat this perplexity as a proxy for adaptation dynamics rather than a calibrated held-out metric, and leave non-overlapping held-out perplexity to future work. Nonetheless, the results show that all three models start with high training perplexity, consistent with a distribution gap between their pre-training mix and Québec French, and drop sharply after the first epoch, indicating rapid uptake of dialectal patterns under CPT.
Figure 1: Perplexity during CPT training. Lower perplexity indicates better fit to Québec French. All models are evaluated on identical held-out corpus.
The results in Table 4 show that all three models improve their performances in Québec French tasks after 6 epochs of CPT. However, these improvements are not uniform across tasks nor models. Out of all four tasks, QFrCoLA proves to be the most difficult one, and most models actually worsen on it or show very small gains after CPT. This may be because of a conflict between the task’s objectives and the training data we used. Indeed, the task consists in labeling a sentence as grammatical or not according to normative Québec French rules; the ungrammatical sentences contain linguistic mistakes that are common in Québec. On the other hand, much of our Québec training data (40%) comes from unedited sources, such as web forum posts and online comments, where these mistakes would be frequently found. The models have thus been trained to accept these erroneous phrasings as “correct”. This would warrant future investigation by conducting CPT exclusively on more formal Québec French sources.
We can note as well that there is a second normative Québec French task, QFrBLiMP, on which our models actually improve after CPT. The difference is that this task is a binary comparison, where the model is given two sentences and must identify the correct from the incorrect one. Taken together, these results indicate that CPT is making the models better at understanding both normative and slang Québec French: the models perform better at distinguishing normative from incorrect language, but also more accepting of incorrect popular language.
Training time seems correlated with performance, as all three models show more improvement after 6 epochs of CPT than 3. However, the correlation to model size is less clear. Llama-3.1-8B clearly outperforms Llama-3.2-1B, but is in turn outperformed by CroissantLLM with 1.35B parameters.
6.2.  Prestige French Retention
As shown in Table 5, two out of the three models, CroissantLLM and Llama-3.1-8B, not only retained their performances on the prestige French tasks but actually improved their average performance, showing that training on a regional dialect helps in handling the prestige dialect as well. Only the smallest model, Llama-3.2-1B, seems to be forgetting general French skills with CPT, and actually becomes worse with more epochs.
The Fr-BoolQ task is the only one whose performance suffers from CPT with all three models. This is probably because our training dataset has no question-answering sources. This, combined with the fact CPT adapts models without replay or regularization, seems to have progressively overwritten the specialized skills needed for this task. However, the other tasks, paraphrase recognition and sentiment detection, require more general language-understanding skills and thus benefit from the expanded language understanding that CPT with Québec French provides (at least for the two models that benefit from CPT).
6.3.  Adaptation-Retention Trade-off
The results in Tables 4 and 5 show that the smaller models, CroissantLLM and Llama-3.2-1B, have difficulty balancing adapting to a new dialect and retaining the skills they have learned previously. CroissantLLM does improve in both Québec French understanding and prestige French ability, but the gains are modest after 3 epochs and the greater gains in the former after 6 epochs almost completely wipe out the gains in the latter. Meanwhile, Llama-3.2-1B’s performance degrades overall throughout the experiment. On the other hand, Llama-3.1-8B, our largest model, shows clear improvements in both Québec French understanding and prestige French ability, and these benefits increase with the number of training epochs. This demonstrates that there is value to using CPT to train models for regional dialects, but only if the base model is large enough to absorb the new information without losing the knowledge gained from its initial training.
7.  Conclusion
In this paper, we demonstrated a case study of low-resource dialectal adaptation of LLMs on Québec French. Our results on 8 tasks from the French benchmark suite COLE show that CPT with LoRA achieves substantial dialectal gains and also improves the model performance on prestige French tasks, but only if the model is large enough to support it. Our work also demonstrated that this specialization can be made by updating $\leq\!1\%$ of model parameters and with using a very small corpus (85M tokens), but that corpus composition has an important impact. Indeed, our corpus being rich in informal and unedited sources, and lacking question-answering sources made the trained models less proficient at distinguishing normatively correct and incorrect text and at question-answering.
As future work, we could explore varying the mix of data sources and examine their influence on performance across specific benchmarks. Further research could also investigate cross-dialect and cross-language CPT, including scenarios that involve code-switching or mixed registers. Finally, exploring techniques such as selective parameter freezing may offer a way to enhance language retention.
8.  Societal Impact and Ethical Considerations
Linguistic equity and preservation.
Unlike general LLMs which are focused on more popular high-resource languages, our work is designed to tailor LLMs to low-resource regional dialects such as Québec French, providing more equitable access to AI tools.
Representation biases and stereotype risks.
Training on naturally occurring Québec slang can strenghten stereotypes linking dialectal features to class, region, or demographics. Our corpus spans news, social media, and forums to diversify exposure, but skews toward written, urban, younger, internet-active speakers while under-representing rural communities, older generations, Indigenous speakers, immigrants, and spontaneous speech. Model behavior may not generalize across these groups. Claims about “Québec French” should be read as conditional on training data demographics.
Dialect subordination via “correction.”
Systems defaulting to prestige French as “correct” (grammar checkers, translation, autocomplete) can normalize away regional features and reinforce linguistic hierarchies. We advocate designs that preserve dialectal variation and distinguish appropriateness from error.
Acknowledgments
This project was undertaken thanks to funding from NSERC, IVADO and the Canada First Research Excellence Fund. The authors would also like to thank CN2i / Le Soleil for generously giving us access to their data.
9.  Limitations
Benchmark coverage and task selection.
Our evaluation focuses exclusively on a subset of the COLE benchmark’s language tasks, which emphasize grammatical acceptability and linguistic structure. This scope excludes sociolinguistic variation (register, formality), conversational competence (authentic dialogue, code-switching), generation quality (dialectal fluency, idiom production), and domain-specific tasks (legal, medical terminology). Human evaluation of generated text would provide critical insights into dialectal authenticity that our automated metrics cannot capture.
Ablation studies and design choices.
We do not report systematic ablations on key hyperparameters: LoRA rank sensitivity (r=16r{=}16 was used following common practice without exploring further values), layer targeting (we apply LoRA to all attention and FFN layers without testing attention-only or selective configurations), gradient checkpointing impact, or adaptive versus uniform sampling strategies. These ablations would strengthen support for our design decisions but were omitted due to computational constraints.
Computational constraints.
Our experiments use modest hardware (single/dual V100 GPUs) with 1024-token sequences, potentially constraining discourse-level phenomena, hyperparameter exploration, and training beyond 6 epochs. While our resource-efficient focus enables broader accessibility, it limits the performance ceiling we can explore.
Access to Québec French data.
High-quality Québec French text is hard to find, scattered, and often behind legal or platform restrictions. Licenses and Terms of Service limit what we can share (e.g., news APIs, social-media comments), and speech transcripts are small. Informal writing exists but is uneven across topics, regions, and age groups. These limits keep our CPT small, make full reproducibility difficult (some sources cannot be released), and may bias models toward urban, online, younger speakers.
10.  Bibliographical References
[…]
Appendix A Training Configuration
Hardware.
  • Devices ($d$): NVIDIA Tesla V100 (32 GB) (single- or multi-GPU as noted below)
  • Precision: fp16 with gradient checkpointing
Batching and Effective Batch. Let $b$ be sequences per device (micro-batch), $a$ the gradient-accumulation steps, and $d$ the number of devices. The effective batch in sequences is
$\text{EffBatch}_{\text{seq}}=b\times a\times d.$
With a sequence length of 1024 tokens, the effective batch in tokens is $\text{EffBatch}_{\text{tok}}=\text{EffBatch}_{\text{seq}}\times 1024$.
Profiles used in our runs.
  • Profile A (1B CPT, single V100): b=4b{=}4, a=8a{=}8, d=1d{=}1 $\Rightarrow$ $\text{EffBatch}_{\text{seq}}{=}32$ and $\text{EffBatch}_{\text{tok}}{=}32{,}768$.
  • Profile B (8B CPT, two V100s): b=1b{=}1, a=8a{=}8, d=2d{=}2 $\Rightarrow$ $\text{EffBatch}_{\text{seq}}{=}16$ and $\text{EffBatch}_{\text{tok}}{=}16{,}384$.
(For larger models, we keep a=8a{=}8 and adjust $b$ and $d$ to fit memory.)
Optimization.
  • Objective: causal language modeling
  • Optimizer: AdamW (weight decay $0.01$)
  • LR schedule: cosine; base LR $1{\times}10^{-5}$; warmup ratio $0.1$
  • Gradient clipping: $1.0$
  • Checkpoint selection: best validation loss per epoch (no early stopping)
LoRA (PEFT).
  • Targets: attention (qkvoqkvo) and FFN ($\mathrm{up,gate,down}$)
  • Rank r=16r{=}16, $\alpha{=}32$, dropout $0.1$
Appendix B LoRA Hyperparameter Ablation
Our main experiments use LoRA with rank r=16r{=}16, $\alpha{=}32$, and dropout $0.1$, following common practice. To validate these choices and provide guidance for future dialect adaptation work, we conducted a systematic ablation study on CroissantLLMChat-v0.1 using our Québec French corpus.
B.1.  Experimental Setup
We evaluated four hyperparameter dimensions: (1) LoRA rank $r\in\{4,8,16,32,64\}$ with $\alpha=2r$; (2) alpha-to-rank ratio $\alpha/r\in\{1,2,4\}$ with fixed r=16r{=}16; (3) dropout $d\in\{0.0,0.05,0.1,0.2\}$; and (4) target modules: $qv$ (query and value projections only), qkvoqkvo (all attention projections), and $Full$ (attention + MLP layers). Each configuration was trained for 3 epochs with identical optimization settings. We report minimum validation loss and perplexity (PPL) achieved during training.
B.2.  Results
Table 6 summarizes the ablation results. We highlight the best configuration within each ablation type.
Ablation Config $r$ $\alpha$ Loss PPL
Rank $r=4$ 4 8 3.057 21.27
$r=8$ 8 16 3.021 20.52
$r=16$ 16 32 2.980 19.68
$r=32$ 32 64 2.930 18.74
$r=64$ 64 128 2.870 17.63
$\alpha/r$ ratioratio $\alpha/r=1$ 16 16 3.009 20.26
$\alpha/r=2$ 16 32 2.980 19.68
$\alpha/r=4$ 16 64 2.953 19.15
Dropout $d=0.00$ 16 32 2.976 19.62
$d=0.05$ 16 32 2.980 19.68
$d=0.10$ 16 32 2.983 19.74
$d=0.20$ 16 32 2.989 19.86
Modules qvqv only 16 32 3.111 22.44
qkvoqkvo 16 32 3.072 21.59
FullFull 16 32 2.979 19.68
Table 6: LoRA hyperparameter ablation on CroissantLLMChat-v0.1 with 3 epochs of CPT on Québec French. Best result per ablation type in bold.
B.3.  Analysis
Rank.
Higher rank consistently improves performance, with r=64r{=}64 achieving $17\%$ lower perplexity than r=4r{=}4 ($17.63$ vs. $21.27$). This suggests that dialect adaptation benefits from increased adapter capacity, likely because capturing lexical, morphological, and syntactic variations requires more expressive low-rank subspaces. However, trainable parameters scale linearly with rank. Examining the efficiency of each doubling: moving from r=32r{=}32 to r=64r{=}64 doubles the parameter count but yields only a $5.9\%$ additional perplexity reduction ($18.74$ to $17.63$), whereas r=32r{=}32 already captures $70\%$ of the total improvement over r=4r{=}4 (perplexity $18.74$ vs. $21.27$). For resource-constrained settings typical of low-resource dialect work, r=32r{=}32 offers a favorable trade-off: near-optimal adaptation quality at half the parameter budget of r=64r{=}64.
Alpha scaling.
The $\alpha/r$ ratioratio controls the effective learning rate of the LoRA updates. We find that $\alpha/r{=}4$ slightly outperforms the standard $\alpha/r{=}2$ setting (loss $2.953$ vs. $2.980$), suggesting that dialect adaptation benefits from more aggressive updates to the low-rank matrices. This may reflect the need to provide stronger gradient updates to shift the model’s lexical and syntactic distributions when adapting to a regional variety.
Dropout.
Contrary to the common practice of using dropout during fine-tuning, we find that zero dropout yields the best results for dialect CPT (loss $2.976$ vs. $2.989$ at d=0.2d{=}0.2). Performance degrades monotonically as dropout increases. We hypothesize that for continual pretraining on a focused domain corpus, the model benefits from fully absorbing dialect-specific vocabulary and expressions. Dropout may interfere with learning these low-frequency but important patterns.
Target modules.
Applying LoRA to all linear layers ($Full$) substantially outperforms attention-only configurations. This configuration achieves $12\%$ lower perplexity than qvqv ($19.68$ vs. $22.44$), indicating that MLP layers play an important role in encoding dialect-specific lexical and semantic knowledge, not just the attention mechanisms.
B.4.  Recommendations
Based on these ablations, the optimal LoRA configuration for low-resource dialect CPT is: $r\geq 32$, $\alpha/r\geq 2$, dropout =0=0, and full module targeting. Our main experiments used r=16r{=}16, $\alpha{=}32$, and d=0.1d{=}0.1 as a conservative baseline following prior work; the ablations suggest that more aggressive settings may yield further improvements. We leave exploration of these optimized settings across all models to future work.